As you stand on a pier looking at a harbor, the waves created by a large freighter come ashore. You estimate the wavelength to be 3.0 m and you count twelve wave peaks arriving in one minute. (a) What is the frequency of the wave in hertz? (b) How fast is the wave traveling in meters per second(m/s)?
1
Expert's answer
2011-04-06T13:26:55-0400
(a) There are 12 wave peaks per 60 seconds: f = 12/60 = 1/5 s-1.
(b) using the following equality λ/v = 1/f, we can find the velocity: v = λf = 3[m] * 1/5 [s-1] = 0.6 m/s.